OpenStudy (anonymous):

H6P1: THE NEWFET DEVICE all questions

OpenStudy (anonymous):

PLZZ help on the H6P2

OpenStudy (anonymous):

1) Minimum bias voltage should be the threshold and the swing Dv; VIN = VT+Dv 2) Here, iDS1=iDS2; 1/2K(VIN-VT)^2=1/2K(VGS2-VT)^2 Therefore, VDS2=VGS2=VIN We know that VDD=VDS1+VDS2 Minimum VDS1(=VOUT) = VIN-VT Therefore, minimum VDD= VIN-VT+VIN =2*VIN-VT 3) gm1= diDS1/dVGS1= K*(VIN-VT) 4) same current, same bias; VGS2=VGS1=VIN gm2= diDS1/dVGS1= K*(VIN-VT) 5) -vin 6) Rth = 1/gm=1/(K*(VIN-VT))

OpenStudy (anonymous):

1) Dv; VIN = VT+Dv 2) 2*VIN-VT 3) gm1= diDS1/dVGS1= K*(VIN-VT) 4) K*(VIN-VT) 5) -vin 6) 1/(K*(VIN-VT)

OpenStudy (anonymous):

@Loveboy54@sam747 i'm so gratful thnk's lot

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

hi can any one help me on question 5 and 6 on the midterm exam